### v1

# FairJudge: MLLM Judging for Social-Attributes and Prompt–Image Alignment

Abstract
Text-to-image (T2I) systems lack simple, reproducible ways to evaluate prompt-image alignment (the core of model accuracy) and social attributes (key to understanding model bias). Standard proxies—face-centric classifiers and contrastive similarity—reward surface correlations, lack calibrated abstention, and have poor coverage of attributes only weakly visible in pixels (e.g., religion, culture, disability). We introduce FairJudge, a lightweight, calibrated protocol that treats instruction-following multimodal LLMs as fair judges for T2I: alignment is scored with an explanation-oriented rubric mapped to [1,1][-1,1], while the judge is constrained to a closed label set, required to ground decisions in explicit visual evidence, and instructed to abstain when cues are insufficient. Unlike the commonly-used CLIP-only pipelines, FairJudge elicits accountable, evidence-aware judgments, and—unlike mitigation methods that intervene on generators—targets evaluation fairness. We evaluate demographics (gender, race, age) on FairFace, PaTA, and FairCoT; extend beyond face-visible cues to religion, culture, and disability; and assess profession correctness and alignment on IdenProf, FairCoT-Professions, and our new DIVERSIFY-Professions. We also contribute DIVERSIFY, a 469-image corpus that varies gender, race, age, religion, culture, and disability under diverse, non-iconic scenes to stress-test evidence-based judging. Across datasets, judge models substantially outperform contrastive and face-centric baselines on demographic prediction and raise mean alignment while maintaining high profession accuracy, yielding a more accurate and transparent pathway for fairness evaluation in T2I.Upon acceptance, we will release all datasets, prompt templates, and code, along with per-image judge outputs, rationales, and parser logs, to enable reproducible audits and re-scoring under alternative rubrics.
Keywords: multimodal LLMs, fairness, text-to-image, social attribute prediction, text–image alignment, evaluation benchmarks
[…]
FairJudge: MLLM Judging for Social-Attributes and Prompt–Image Alignment
Zahraa Al Sahili, Maryam Fetanat*thanks: *Equal contribution, Maimuna Nowaz* 11footnotemark: 1, Ioannis Patras, Matthew Purver
Queen Mary University of London, Imperial College London, Institut Jožef Stefan
London, UK, London, UK, Ljubljana, Slovenia
maryam.fetanat20@imperial.ac.uk, m.nowaz@se23.qmul.ac.uk
{z.alsahili, i.patras, m.purver}@qmul.ac.uk
Abstract content
1.  Introduction
Text-to-image (T2I) systems are increasingly used to generate and retrieve people-centric imagery in creative, assistive, and analytical settings. Yet these models inherit—and can amplify—social biases from their data and training objectives, yielding skewed portrayals across gender, skin tone, geo-culture, religion, disability, and professions. Conventional evaluation pipelines rely heavily on face-centric attribute recognizers or contrastive similarity scores to stand in for human judgment. Such proxies struggle when the target attribute is weakly encoded in pixels (e.g., culture), when confounding context is present (e.g., clothing, scene props), or when the “right” decision is to abstain. As a result, fairness assessments may either overstate model capability or mask failure modes that would be obvious to a careful human reviewer.
Large-scale studies have documented representation disparities in T2I and vision–language models and have introduced diagnostic tests and scorecards for auditing them (Luccioni et al., 2023; Wan et al., 2024). Remedial approaches range from lightweight sampling control to small learned adaptors. Fair Diffusion steers the sampler toward user-specified demographic proportions without retraining the backbone (Friedrich et al., 2023); Fair Mapping inserts a linear adapter to systematically reduce demographic skew with minimal compute (Li et al., 2023). Generation-time prompting with chain-of-thought (CoT), as in FairCoT, leverages reasoning to mitigate bias during image synthesis (Al Sahili et al., 2024). While effective for mitigation, these approaches operate on the generator; they do not provide an evaluation protocol that is calibrated, evidence-grounded, and appropriately abstaining. On the evaluation side, contrastive similarity (e.g., CLIP) is a dominant proxy for text–image faithfulness (Radford et al., 2021; Hessel et al., 2021), and VQA-style checklists probe finer-grained grounding (Hu et al., 2023), but both can reward surface correlation and penalize uncertainty. Preference-trained rankers (ImageReward, PickScore) improve agreement with human choices (Xu et al., 2023; Kirstain et al., 2023) yet remain scalar scorers with limited transparency. Recent “LLM-as-a-judge” work reports high agreement with human raters but also reveals position and verbosity biases, and thus calls for careful protocol design and meta-evaluation (Zheng et al., 2023; Chen et al., 2024; Wang et al., 2025). In short: we lack a lightweight, evaluation-oriented procedure that (i) reasons explicitly about visual evidence, (ii) adheres to label constraints, (iii) abstains when cues are insufficient, and (iv) applies beyond face-visible demographics to attributes like religion, culture, disability, and profession.
Figure 1: CLIP is brittle for profession recognition; MLLM judges are consistent. Each panel shows the same profession under appearance/cultural variation.
To address these gaps, we propose FairJudge, a calibrated protocol that treats instruction-following multimodal LLMs (MLLMs) as fair judges for T2I evaluation. Technically, FairJudge (1) enforces label-constrained outputs for target attributes and professions; (2) requires the judge to cite explicit, observable visual evidence and to return a neutral/unknown outcome when evidence is missing; (3) standardizes alignment scoring with an explanation-oriented rubric that maps to a continuous [1,1][-1,1] scale; and In contrast to CLIP-only pipelines that score similarity, FairJudge elicits accountable, evidence-aware decisions; and unlike generation-time debiasing (FairDiffusion, FairMapping, FairCoT), our focus is evaluation fairness.
Contributions.
(i) We introduce FairJudge, a lightweight, label-constrained, abstention-capable judging protocol for social-attribute prediction and prompt–image alignment. (ii) We contribute two new resources—DIVERSIFY and DIVERSIFY-Professions—that probe culturally situated attributes and occupational grounding beyond face-visible cues. (iii) We provide a comprehensive evaluation across FairFace, PaTA, FairCoT, DIVERSIFY (gender, race, age, religion, culture, disability) and IdenProf, FairCoT-Professions, DIVERSIFY-Professions (alignment), showing consistent gains over CLIP and DeepFace, with the largest margins where contextual reasoning matters. (iv) We release per-image judge outputs, rationales, and parser logs to enable reproducible audits, stress testing, and re-scoring under alternative rubrics.
Taken together, our results indicate that calibrated MLLM judges close long-standing gaps in T2I evaluation—capturing non-facial, culturally grounded attributes and producing interpretable alignment scores—while remaining easy to implement and reproducible across rapidly evolving models. We release code, prompts, and parsers for replication upon acceptance.
2.  Related Work
2.1.  Bias in Text-to-Image Models
Text-to-image (T2I) systems inherit and sometimes amplify social biases present in their training data. Early large-scale analyses show systematic disparities across gender, skin tone, geo-culture, and professions, and propose diagnostic protocols and scorecards for auditing representation (Luccioni et al., 2023; Wan et al., 2024). Several debiasing methods have been explored post-hoc or with light adaptation: Fair Diffusion steers sampling toward user-specified demographic proportions (Friedrich et al., 2023), while Fair Mapping learns a small linear adaptor to systematically reduce demographic skew with minimal compute (Li et al., 2023). Concurrently, FairCoT leverages chain-of-thought (CoT) reasoning in multimodal LLMs to mitigate bias during generation-time prompting (Al Sahili et al., 2024). Our work differs in purpose: rather than intervening on the generator, we study evaluation fairness by treating multimodal LLMs (MLLMs) as judges over social attributes and alignment outcomes.
$2.$  Evaluating T2I: Attribute Prediction vs. Alignment
Attribute prediction.
Face-centric attribute benchmarks remain widely used to quantify gender, race, and age performance. FairFace provides a balanced dataset across seven race groups and age/gender labels (Kärkkäinen and Joo, 2021). For context-embedded protected-attribute cues beyond faces, the PaTA benchmark was curated alongside the DeAR debiasing method for VLMs (Seth et al., 2023). Tooling spans classic face-attribute models and modern vision-language recognition. In practice, many works rely on the open-source DeepFace toolkit to obtain age, gender, and race predictions (Serengil, 2024), and on CLIP (Radford et al., 2021) for zero-shot category attribution via text prompts. However, audits of CLIP-family encoders find systematic gender and racial skew whose magnitude depends on model size and, crucially, the pre-training data source/scale (Al Sahili et al., 2025). We extend this line by prompting MLLMs (GPT-4.1, Gemini, LLaMA-4) to directly infer social attributes, enabling coverage of non-facial categories (e.g., religion, culture, disability) with explanations.
Text–image alignment.
CLIP-based similarity remains a dominant automatic proxy for prompt faithfulness (Hessel et al., 2021). Beyond cosine-based scores, VQA-driven faithfulness frameworks such as TIFA generate question–answer checklists to probe fine-grained concept grounding (Hu et al., 2023). Preference-trained reward/score models (e.g., ImageReward and PickScore) leverage large-scale human comparisons to better match human judgments for T2I ranking and evaluation (Xu et al., 2023; Kirstain et al., 2023). Recently, X-IQE proposes a Chain-of-Thought, explanation-oriented evaluator using visual LLMs, reporting improved agreement with humans while outputting interpretable rationales (Chen et al., 2023). In our alignment study, we compare CLIP similarity against MLLM-based judging calibrated with an X-IQE–style scoring protocol to a standardized [1,1][-1,1] range.
2.3.  LLMs/MLLMs as Judges
Using LLMs as evaluators (“LLM-as-a-judge”) has become a practical alternative to costly human evaluation. Foundational work reports high agreement with human preferences but also identifies position, verbosity, and self-enhancement biases, motivating careful protocol design and meta-evaluation (Zheng et al., 2023). Subsequent studies investigate reliability, consistency, and bias in LLM judging, including on open-ended tasks (Chen et al., 2024). For images, recent work shows that MLLMs can serve as safety and policy judges with structured reasoning pipelines, revealing both promise and pitfalls for multimodal evaluation (Wang et al., 2025). Our work complements this literature by: (i) framing social-attribute prediction as a judging task on face and non-face domains (e.g., culture, religion, disability), (ii) quantifying per-attribute accuracy and generalization across FairFace, PaTA, and new human-centric scenarios, and (iii) benchmarking alignment judging on IdenProf professions and FairCoT—contrasting CLIP-based scores with MLLM judges that also supply free-form rationales.
3.  Methods
3.1.  Background: CLIP and CLIP-based Alignment
CLIP. Contrastive Language–Image Pretraining (CLIP) learns an image encoder $f_{\theta}(\cdot)$ and a text encoder $g_{\phi}(\cdot)$ such that paired image–text embeddings have high cosine similarity while mismatched pairs are pushed apart. Given an image $I$ and a textual prompt $t$, the canonical alignment score is the cosine
$\mathrm{CLIPAlign}(I,t)=\cos\!\big(f_{\theta}(I),\,g_{\phi}(t)\big)\in[-1,1].$
For multi-class recognition with a label set $\mathcal{Y}$, one typically forms templated prompts $\{t_{y}:y\in\mathcal{Y}\}$ and predicts $\arg\max_{y\in\mathcal{Y}}\mathrm{CLIPAlign}(I,t_{y})$, optionally temperature-scaled into a softmax over classes.
Limitations for fairness evaluation. CLIP-based scores are sensitive to prompt wording and class name choices; they provide no calibrated abstention when evidence is weak; and they reward surface correlations, especially for attributes only partially encoded in pixels (e.g., culture, religion, disability). In addition, CLIP cannot natively justify decisions or enforce label constraints beyond the provided prompts, which limits its utility as an auditable evaluator.
3.2.  MLLM-as-Judge Protocol
We treat instruction-following multimodal LLMs (MLLMs) as judges. The protocol is lightweight and calibrated: (i) label constraints restrict outputs to a closed set (with unknown/unspecified as a neutral fallback); (ii) the model must cite explicit visual evidence for each decision; and (iii) the judge abstains when evidence is insufficient. We instantiate judges with GPT-4.1, Gemini 1.5 pro, and LLaMA-4; prompts are identical across models.
3.2.1.  Social-Attribute Prediction
We frame attribute recognition as a single, structured judgment over six targets: gender, race, age, religion, culture, and disability. The judge receives the image and an instruction that (a) enumerates admissible labels (including unknown/unspecified), (b) requests a strict JSON output with fixed fields, and (c) requires a short, image-grounded evidence rationale per attribute. For each model, we instantiate the same schema and constraints with minor phrasing adjustments (e.g., synonyms, formatting reminders) that do not alter the decision rules. An abridged template appears below; complete prompts will be provided upon acceptance.
You are given an image. Predict the following attributes of the person in the image:
1. Gender (male/female/unspecified)
2. Race (Black, White, Asian, Latino_Hispanic, Indigenous, unspecified)
3. Age (child, young adult, middle-aged, elderly, unspecified)
4. Religion (Christian, Muslim, Hindu, Buddhist, Jewish, Sikh, Shinto, Neutral, unspecified)
5. Culture (e.g., British, African, East Asian, Latin American, or unspecified)
6. Disability (mobility impairment, blind/low vision, deaf/hard of hearing, dwarfism, vitiligo, unspecified)
Output strict JSON with keys: "Gender", "Race", "Age", "Religion", "Culture", "Disability". If evidence is insufficient, use "unspecified".
Calibration and neutrality. We adopt two safeguards. First, we insert explicit instructions to only select labels supported by visible cues; otherwise output unspecified. Second, we request a terse rationale per field during development to audit evidence use; at evaluation time, we parse the JSON keys only (rationales are logged for reproducibility and spot checks). This design reduces overconfident guessing and makes abstentions systematic rather than ad hoc.
3.2.2.  Prompt–Image Alignment Judging
We score prompt–image alignment using a rubric-based judgment inspired by explanation-oriented evaluators (e.g., X-IQE). Given image $I$ and description $t$, the judge assigns a discrete rating $r\in\{1,\dots,5\}$ using the following anchors:
According to the image and your previous description, how well does the image align with the following description?
(1) Not match at all; (2) Has significant discrepancies; (3) Several minor discrepancies; (4) A few minor discrepancies; (5) Matches exactly.
We then map $r$ to a continuous alignment score $\overline{s}\in[-1,1]$ via a linear rescaling,
$\overline{s}=\frac{r-3}{2},\qquad\text{so that }1\mapsto-1,\;3\mapsto 0,\;5\mapsto 1,$
making it directly comparable to cosine-based CLIP alignment. As with attributes, the prompt requires evidence-backed justifications during development; at evaluation we use the scalar $\overline{s}$ while retaining rationales for audit. To assess instruction sensitivity, we evaluate with minimal template perturbations (e.g., “a photo of” $\leftrightarrow$ “an image of”) and confirm that model ordering is stable even if absolute values move.
(a) Attribute prediction (MLLM as judge). The image and label-constrained prompt are fed to the judge, which outputs strict JSON for Gender, Race, Age, Religion, Culture, and Disability, with calibrated abstention.
(b) Prompt–image alignment (MLLM as judge). Given an image and caption, the judge assigns a rubric score $r\!\in\!\{1,\dots,5\}$ with a brief justification; we rescale to $\overline{s}\!\in[-1,1]$ for direct comparison with CLIP cosine.
Figure 2: FairJudge: instruction-following MLLMs as fair judges. (a) Social-attribute prediction with label constraints and abstention; (b) rubric-based alignment compatible with CLIP’s scale. Rationales are logged for transparency but not used for scoring.
$3.$  Dataset Contributions: DIVERSIFY and DIVERSIFY-Professions
We introduce two complementary resources that extend fairness evaluation beyond face-visible cues and iconic settings.
DIVERSIFY.
DIVERSIFY comprises 469 synthetic images rendered from controlled prompts and paired with normalized metadata for six targets: gender, race, age, religion, culture/ethnicity, and disability. The corpus is designed to surface context-dependent, culturally situated evidence (attire, objects, signage, scene) where simple face heuristics are insufficient. Demographic coverage is balanced by design (gender $\approx$ 50/50), with race and age distributed across meaningful strata (e.g., Black/White/East Asian/Middle Eastern/Latino_Hispanic; young-adult/middle-aged/elderly). Beyond demographics, DIVERSIFY explicitly varies religion (including Neutral where no visible symbol appears), culture/ethnicity (single tag in this release, spanning Latin American, South/East Asian, African, Western/British, etc.), and disability (mobility aids, blind/low vision, deaf/hard of hearing, dwarfism, vitiligo, among others, with unspecified when absent). Prompts are composed as base scenes (indoor/outdoor; public/private; formal/informal) with single-factor variants that toggle exactly one attribute while holding others fixed, enabling counterfactual probes. Labels are text-derived from the prompt taxonomy; we apply deterministic parsing, canonicalization (e.g., Indian $\rightarrow$ South Asian), near-duplicate filtering, and sanity checks for symbol/object presence.
DIVERSIFY-Professions.
To evaluate occupational grounding and alignment under the same diversity principles, DIVERSIFY-Professions contains 1200 images organized as 6 professions $\times$ 200 (doctor, engineer, janitor, lawyer, nurse, teacher). Each example pairs an image with a profession description and is constructed in non-iconic, culturally varied scenes to reduce shortcut cues. We report (i) top-1 profession accuracy and (ii) a rubric-based alignment score $\overline{s}\in[-1,1]$ derived from a 1–5 judge rating. Distractor professions are selected to be visually plausible, increasing the need for explicit, evidence-backed decisions. Metadata mirror DIVERSIFY (gender, race, age, religion, culture cues), enabling slice-wise fairness analysis conditioned on profession.
Summary. Together, DIVERSIFY (469 images; six target attributes including religion, culture, and disability) and DIVERSIFY-Professions (1,200 images; six professions) furnish controlled, culturally aware evaluations that complement FairFace/PaTA/FairCoT by exposing context-driven failure modes and enabling counterfactual, slice-specific audits.
Figure 3: Overview of the DIVERSIFY benchmark. A mosaic of example images illustrating diverse professions, cultures, scenes, and non-iconic viewpoints. The dataset is designed to reduce shortcut cues and stress-test judges on religion, culture, and disability situated, context-dependent signals.
4.  Experimental Results
We evaluate three questions: (i) social-attribute prediction for gender, race, and age on FairFace, PaTA, FairCoT, and DIVERSIFY; (ii) generalization to religion, culture, and disability; and (iii) prompt–image alignment on IdenProf, FairCoT-Professions, and DIVERSIFY-Professions. For alignment we report top-1 profession accuracy and mean alignment $\overline{s}\!\in[-1,1]$.
4.1.  Experimental Setup
Datasets.
For attribute prediction we use FairFace (balanced faces with gender/race/age), PaTA (protected attributes embedded in broader scenes), FairCoT (faces with additional non-facial context), and DIVERSIFY (images curated to surface culture, religion, disability alongside demographics). For alignment we use IdenProf (single-subject professions), FairCoT-Professions (profession cues with face/context), and DIVERSIFY-Professions (non-iconic, culturally diverse settings).
Models and baselines.
Judges are GPT-4.1, Gemini 1.5 pro, and LLaMA-4, prompted with identical, label-constrained instructions (Sec. 3.2). Baselines are CLIP (zero-shot classification via templated prompts; cosine similarity for alignment) and DeepFace (face-centric gender/race/age).
Metrics.
For attributes we report top-1 accuracy per attribute and dataset, plus macro means across attributes/datasets (populated cells only). For alignment, judges assign a 1–5 rubric score which we linearly rescale to $\overline{s}\in[-1,1]$ and we also report profession accuracy where annotated. We include minimal template perturbations (“a photo of”/“an image of”) to check instruction sensitivity.
4.2.  Social-Attribute Prediction
Overall trends.
Across all datasets, judge models deliver consistent, double-digit improvements over CLIP and clearly exceed DeepFace on gender and age. Race remains the most challenging attribute, but judges still narrow the gap relative to contrastive and face-centric baselines. The pattern holds across both face-centric (FairFace, FairCoT) and context-rich (PaTA, DIVERSIFY) settings, indicating that evidence-seeking instructions help even when cues are subtle or distributed across the scene.
By dataset.
On FairFace, GPT-4.1 leads on race and age, while LLaMA-4 slightly surpasses others on gender; CLIP is competitive on gender but lags notably on age. On PaTA, GPT-4.1 attains the strongest race and age performance, with Gemini 1.5 pro and LLaMA-4 tying for the top gender score. On FairCoT, Gemini 1.5 pro tops gender, GPT-4.1 tops race, and LLaMA-4 tops age, reflecting complementary strengths. On DIVERSIFY, where demographic cues are less iconic, judges still outperform CLIP by large margins on gender and age and offer the strongest overall coverage. Detailed numbers appear in Table 1; best scores per column are bolded within each dataset block.
Table 1: Social-attribute prediction (top-1 accuracy). Columns: Gender, Race, Age.00footnotetext: DeepFace results on FairFace use 70% of the dataset; full-dataset results will be updated for the camera-ready version.
Dataset Method Gender Race Age
FairFace DeepFace 0.750 0.480 0.620
FairFace CLIP 0.940 0.690 0.820
FairFace GPT-4.1 0.960 0.880 0.980
FairFace Gemini 1.5 pro 0.950 0.784 0.741
FairFace LLaMA-4 0.970 0.710 0.820
PaTA CLIP 0.840 0.740 0.480
PaTA DeepFace 0.750 0.520 0.620
PaTA GPT-4.1 0.930 0.920 0.960
PaTA Gemini 1.5 pro 0.990 0.820 0.960
PaTA LLaMA-4 0.990 0.710 0.950
FairCoT CLIP 0.780 0.700 0.910
FairCoT DeepFace 0.520 0.500 0.840
FairCoT GPT-4.1 0.970 0.820 0.930
FairCoT Gemini 1.5 pro 0.990 0.820 0.960
FairCoT LLaMA-4 0.990 0.690 0.970
DIVERSIFY CLIP 0.782 0.328 0.546
DIVERSIFY GPT-4.1 0.981 0.476 0.812
DIVERSIFY Gemini 1.5 pro 0.992 0.463 0.934
DIVERSIFY LLaMA-4 0.981 0.400 0.840
Judge instructions that enforce label constraints, require visible evidence, and allow abstention translate into higher accuracy on attributes anchored in appearance (gender, age) and competitive gains on harder, context-affected targets (race). The leadership swaps across datasets point to complementary strengths of the MLLMs rather than a single dominant model.
4.3.  Generalization to Religion, Culture, and Disability
Judge models extend beyond face-visible demographics. On FairCoTreligion, GPT-4.1 and LLaMA-4 achieve high accuracy while CLIP trails substantially, reflecting the value of evidence-based reasoning for symbols and context. On DIVERSIFY, judges outperform CLIP by large margins on religion and disability; culture remains the hardest category, with smaller gaps—consistent with diffuse, relational cues that are often underdetermined by the pixels. Full results appear in Table 2.
Table 2: Generalization beyond demographics (accuracy).
Dataset Method Religion Culture Disability
DIVERSIFY CLIP 0.345 0.322 0.360
DIVERSIFY GPT-4.1 0.689 0.258 0.932
DIVERSIFY Gemini 1.5 pro 0.817 0.271 0.795
DIVERSIFY LLaMA-4 0.700 0.397 0.902
FairCoT CLIP 0.410
FairCoT GPT-4.1 0.880
FairCoT Gemini 1.5 pro 0.820
FairCoT LLaMA-4 0.880
Where attributes require contextual reasoning or symbol recognition, judges offer a clear advantage over contrastive similarity. Culture remains an open challenge, reinforcing the importance of calibrated abstention and dataset design that minimizes proxy shortcuts.
$4.$  Prompt–Image Alignment
Judges also improve semantic grounding between descriptions and images. On IdenProf, judge models match or exceed CLIP on profession accuracy and substantially increase the mean alignment score $\overline{s}$. On FairCoT-Professions, GPT-4.1 yields the highest alignment with competitive accuracy, and on DIVERSIFY-Professions all judges surpass CLIP by a wide margin on alignment while maintaining strong accuracy. Minimal template changes do not alter model ordering, indicating robustness of the judging rubric. Table 3 reports full results.
Table 3: Prompt–image alignment. We report top-1 Accuracy and mean Alignment ([1,1][-1,1]).
Dataset Method Accuracy Alignment
DIVERSIFY-Professions CLIP 0.736 0.250
DIVERSIFY-Professions GPT-4.1 0.862 0.810
DIVERSIFY-Professions Gemini 1.5 pro 0.842 0.770
DIVERSIFY-Professions LLaMA-4 0.676 0.526
FairCoT-Professions CLIP 0.471 0.250
FairCoT-Professions GPT-4.1 0.670 0.710
FairCoT-Professions Gemini 1.5 pro 0.661 0.702
FairCoT-Professions LLaMA-4 0.685 0.448
IdenProf CLIP 0.927 0.270
IdenProf GPT-4.1 0.937 0.680
IdenProf Gemini 1.5 pro 0.943 0.709
IdenProf LLaMA-4 0.888 0.370
Rubric-calibrated judging produces alignment scores that better reflect profession correctness and scene semantics than cosine similarity, while avoiding over-penalization from prompt phrasing.
4.5.  Ablation
We conduct an ablation study on the FairCoT dataset to probe the contribution of each protocol component—label constraints, abstention via unspecified, and evidence prompting. For scoring, we apply an attribute-specific policy for unspecified (abstentions): (i) for gender, race, and age, unspecified is treated as a misclassification (i.e., counted in the denominator with zero credit); (ii) for religion and culture, unspecified is treated as neutral and excluded from accuracy (we also report coverage as the fraction of non-abstentions); and (iii) for disability, unspecified is interpreted as predicting no disability (i.e., scored as correct when the ground truth is “no disability” and as an error otherwise). Table 4 reports per-attribute accuracy under three settings per model: (i) Without labels (free-form answers), (ii) without unspecified (forced prediction), and (iii) our full FairJudge protocol.
Table 4: Ablations on FairCoT (per-attribute accuracy). “Without labels” removes the closed taxonomy; “without unspecified” disables abstention. LLaMA age reflects the author’s normalization; religion notes indicate frequent abstention/free text.
Model Setting Gender Race Age Religion
GPT-4.1 Without labels 94.7% 11.2% 0.86% 59.5%
GPT-4.1 without unspecified
GPT-4.1 our 97% 82% 93% 88%
Gemini 2.5 Pro Without labels 80.17% 17.24% 0% 85.34%
Gemini 2.5 Pro without unspecified 92.20% 15.80% 8.40% 21.10%
Gemini 1.5 Pro our 99% 82% 96% 82%
LLaMA-4 Without labels 43% 21.40% 89% 10%
LLaMA-4 without unspecified 45% 35% 90% 9%
LLaMA-4 our 99% 69% 97% 88%
“(Normalised)” in source sheet.    “Most said cannot be determined” (frequent abstention/free-text).
Removing the label constraints causes steep drops on the hardest axes (race, age): free-form outputs drift off the taxonomy and cannot be reliably parsed back, even when the model’s prose sounds reasonable. Disabling abstention forces guesses and disproportionately harms attributes with weak or intermittent evidence (religion, age), as seen most clearly for Gemini. By contrast, the full protocol—with a closed label set, an explicit unspecified option, and an evidence-seeking instruction—recovers high gender/age accuracy and substantially improves race/religion, especially for LLaMA-4 where the gap between “without labels” and “our” is largest. In short: constraints prevent drift; abstention curbs overconfidence; evidence prompts focus the judge on verifiable cues. These effects are strongest exactly where unconstrained or forced-guess pipelines are most brittle.
4.6.  Computational Cost
We report wall-clock time per image and marginal API cost per image exactly as recorded in our logs.
Table 5: Cost/speed per image
Model Time / image Cost / image
LLaMA-4 0.0789 min \$0.0817
CLIP 0.21 s open-source
GPT-4.1 4.47 s \$0.0025
Gemini 1.5 Pro 14.42 s \$0.0017
CLIP offers unmatched throughput and negligible marginal cost but provides the least calibrated fairness/faithfulness signal for our purposes, so we focus on MLLM judges. Among judges, GPT-4.1 achieves a strong quality–cost trade-off (low per-image cost and moderate latency), while Gemini 1.5 Pro is the cheapest per image but slower in practice. Under our setup, LLaMA-4 is the most costly because we accessed it via the DeepInfra API; if self-hosted as open source, the marginal API cost would be zero but it would be the most computationally expensive to run (GPU-hours/energy/ops). A pragmatic deployment is a progressive judging cascade: route easy cases to a cheaper judge (e.g., Gemini 1.5 Pro), escalate only abstentions or low-evidence cases to GPT-4.1, and invoke LLaMA-4 only for disagreements between judges or fairness-critical audits requiring a second opinion. This preserves most of the accuracy/alignment gains while controlling runtime and spend, and it aligns with an auditor’s workflow where explanations, abstentions, and cross-checks matter most.
Across attributes, non-demographic categories, and alignment, instruction-following judges offer a practical path to accountable evaluation: higher accuracy where cues are visible, better grounding where cues are subtle, and stable rankings under small prompt changes.
5.  Discussion
Our results show that judge-style multimodal LLMs provide a more faithful and conservative read of sensitive attributes than contrastive or face-centric baselines. The strongest gains appear where visual evidence is diffuse or context-dependent. On gender and age, judges reliably outperform CLIP and DeepFace across datasets, and on religion and disability they close large gaps that face-only tools fail to bridge (Tables 1, 2). In prompt–image alignment, judges also raise mean alignment substantially over CLIP-$T$ while maintaining high profession accuracy (Table 3). Together, these findings suggest that judge models are measuring something closer to “task intent with evidence,” rather than surface similarity alone.
A key reason for this improvement is the use of label-constrained prompting and neutral fallbacks. By requiring the model to ground its decisions in observable cues and to abstain otherwise, we reduce overconfident guesses that often plague open-ended evaluators. This shift is most visible in domains like religion and disability, where naive heuristics from facial attributes are inappropriate. The judges’ behavior resembles a cautious annotator: they succeed when there is verifiable evidence in the image and resist filling gaps with speculation. In contrast, contrastive models continue to shine on attributes that are strongly anchored in the pixels and annotation conventions (for example, gender on FairFace), but they do not offer the same calibrated abstention or rationale structure.
The approach is complementary to representation learning. In practical pipelines, contrastive similarity can still propose candidates with high recall, while the judge acts as a verifier that enforces label constraints and discards spurious matches. This division of labor is especially effective in professions and other multi-attribute settings, where retrieving plausible categories is easy but verifying the precise category with defensible evidence is not. Our template perturbation check further indicates that while absolute scores move with phrasing, the judge-based ordering is stable, reinforcing the view that judge outputs capture semantically grounded correspondences rather than brittle lexical artifacts.
Cultural attributes remain challenging. Unlike religion, which often manifests through specific, visually salient symbols that a cautious judge can check for, culture is diffuse, relational, and frequently non-visual. The smaller margins over CLIP on culture are therefore unsurprising and underscore the limits of image-only evaluation. Equally, the per-dataset leaders vary—FairFace rewards certain cues differently than DIVERSIFY—which reminds us that “difficulty” is a property of both the phenomenon and the annotation regime. As a result, cross-dataset calibration and explicit reporting by attribute and dataset are not just good practice but necessary to avoid overclaiming generality.
Finally, there are safety considerations. Judging sensitive attributes can cause harm if labels are ill-defined or spuriously inferred. Our protocol mitigates this by constraining outputs, privileging explicit evidence, and encouraging abstention. In deployment, audits should be run offline to characterize models and datasets rather than people, and decisions that affect individuals should never be automated from these inferences. Within that scope, judge-style evaluation offers a clearer and more accountable lens on fairness than score-only pipelines.
6.  Conclusion
We introduced a lightweight, calibrated protocol that treats instruction-following MLLMs as fair judges for text–to–image (T2I) evaluation. By constraining outputs to label sets, requiring explicit visual evidence, and encouraging principled abstention, the protocol shifts evaluation from score-only heuristics toward accountable, evidence-aware judgments. Empirically, judge models deliver consistent gains over contrastive and face-centric baselines on gender and age, and close large gaps on attributes that demand non-facial reasoning, notably religion and disability. They also increase prompt–image alignment while retaining high profession accuracy, indicating that their scores better track task intent rather than surface correlations.
A second contribution is the release of DIVERSIFY and DIVERSIFY-Professions, which probe fairness and grounding beyond face-visible cues. These resources expose failure modes that demographic benchmarks miss, revealing that culture remains challenging and often underdetermined by pixels alone. The combination of a calibrated judging protocol and culturally oriented testbeds yields a clearer picture of what current systems can—and cannot—reliably infer.
Taken together, our results support a practical design pattern: pair contrastive retrieval for high-recall candidate generation with judge-style verification for precision, justification, and calibrated abstention. This cascaded approach preserves the efficiency of representation learning while inserting an accountable reasoning step where it matters. To foster rigor and transparency, we release per-image judge outputs, rationales, and parser logs alongside the new datasets, enabling independent analysis, stress testing, and re-scoring under alternative rubrics.
Important limitations remain. Cultural labels are fluid and context-dependent; instruction phrasing can shift absolute scores; and domain shift can reorder leaders across datasets. We view these as invitations for broader stress tests, cross-dataset calibration, and triangulation with human evaluation. Nevertheless, the core message is clear: judging with evidence and evaluating beyond faces offer a concrete path toward safer, fairer, and more interpretable assessment of multimodal systems.
7.  Limitations
Despite the gains, several limitations deserve emphasis. First, ambiguity in annotation undermines ceiling performance, particularly for culture where definitions are fluid and culturally situated. Even when judges abstain appropriately, measured accuracy will reflect the dataset’s operationalization more than any universal construct. Second, some attributes may not be visually inferable from a single image. High scores can arise from dataset-specific regularities rather than genuine visual grounding; conversely, responsible abstention may be penalized when labels require extra-visual context. Third, judge outputs remain sensitive to instruction framing. We reduce this with label constraints, neutral defaults, and template checks, but broader stress testing across prompts and task wordings is needed to quantify instruction variance.
Fourth, dataset shifts matter. The identity of the top method changes across datasets and attributes, indicating residual coupling to collection practices and label taxonomies. Robust conclusions therefore require leave-one-dataset-out analyses, calibrated thresholds, and reporting of uncertainty (e.g., bootstrap intervals), which we encourage but have not exhaustively explored here. Fifth, judges can still exploit ethically fraught correlates. Evidence-grounding reduces spurious inference but does not eliminate the possibility that a model leverages proxy cues. Audits should thus include qualitative review of rationales and, where possible, counterfactual probes that break such proxies.
Sixth, the alignment rubric introduces metric dependence. While our rubric correlates with accuracy and improves semantic grounding, different rubrics or panel judgments could shift absolute values. Future studies should triangulate with human evaluation and alternative scoring schemes. Finally, computational cost and latency are higher for judge models than for single-pass contrastive scoring, which may limit their use at scale. We view the added cost as justified for fairness auditing, but practical deployments will benefit from cascaded designs, caching, and selective judging only where contrastive filters indicate uncertainty or sensitivity.
8.  Bibliographical References
[…]
Appendix A Dataset Contribution and Datasheet: DIVERSIFY
A.1.  Overview
DIVERSIFY is a synthetic image corpus for fairness evaluation beyond face-visible cues. Each image is rendered from a text prompt and paired with structured metadata for six targets: gender, race, age group, religion, culture/ethnicity, and disability. Prompts are composed as base scenes (e.g., indoor/outdoor, public/private, formal/informal) with systematic variants that toggle one factor at a time, encouraging controlled analyses rather than opportunistic mining of web photos. The dataset is designed to surface context-dependent, culturally situated evidence (attire, objects, signage, setting) where simple face heuristics are insufficient.
Intended use.
Primary uses include: (i) measuring recognition and abstention behavior of evaluators (CLIP and MLLM judges) on attributes with weak pixel evidence; (ii) auditing spurious correlations between demographics and scene or cultural cues; and (iii) running counterfactual tests by swapping a single attribute while holding the rest constant. The corpus is not intended for training identity recognizers or for inferring attributes of real people.
A.2.  Composition and Statistics
Each example has a unique prompt ID, the rendered image, and normalized labels parsed from prompt text. Demographic attributes (gender, race, age) are single-label; religion is single-label with unspecified; culture/ethnicity is multi-label to reflect multiple cues; disability is single-label with unspecified. Below we provide summary tables mirroring DIVERSIFY-Professions. Counts depend on the specific release split and can be filled from the provided metadata CSV. [^1] We release per-image metadata with canonicalized labels and prompt text; see §A.3.
Table 6: Gender distribution ($N=469$).
Label Count %
Male 234 49.9%
Female 235 50.1%
Unspecified 0 0.0%
Table 7: Race distribution (canonicalized) ($N=469$). Indian $\rightarrow$ South Asian.
Label Count %
Black 109 23.2%
White 167 35.6%
East Asian 93 19.8%
South Asian 4 0.9%
Middle Eastern 45 9.6%
Latino/Hispanic 48 10.2%
Indigenous 0 0.0%
Southeast Asian 3 0.6%
Table 8: Age groups ($N=469$).
Label Count %
Child 0 0.0%
Young adult 157 33.5%
Middle-aged 151 32.2%
Elderly 151 32.2%
Unspecified 10 2.1%
Table 9: Religion distribution ($N=469$).
Label Count %
Neutral 130 27.7%
Christian 47 10.0%
Muslim 54 11.5%
Hindu 46 9.8%
Buddhist 49 10.4%
Jewish 57 12.2%
Sikh 54 11.5%
Shinto 32 6.8%
Unspecified 0 0.0%
Table 10: Culture/Ethnicity tags (single label in this release; $N=469$).
Culture tag Prompts % of NN
Unspecified 180 38.4%
Latin American 60 12.8%
South Asian 57 12.2%
Western 54 11.5%
African 51 10.9%
East Asian 33 7.0%
British 9 1.9%
West African 5 1.1%
Gulf 5 1.1%
East African 3 0.6%
Southern African 3 0.6%
Southeast Asian 3 0.6%
Persian 1 0.2%
Pacific Islander 1 0.2%
Mediterranean 1 0.2%
Central African 1 0.2%
North African 1 0.2%
Eastern European 1 0.2%
Table 11: Disability labels ($N=469$).
Label Count %
Unspecified 409 87.2%
Mobility impairment 17 3.6%
Blind/low vision 10 2.1%
Down syndrome 10 2.1%
Amputee 6 1.3%
Deaf/hard of hearing 4 0.9%
Dwarfism 4 0.9%
Autism 4 0.9%
Vitiligo 2 0.4%
Albinism 2 0.4%
Cerebral palsy 1 0.2%
Label provenance.
All labels are text-derived from the prompt templates; no human annotation of rendered pixels is required. To avoid hallucinated labels, we canonicalize descriptors to a closed taxonomy and map absent evidence to unspecified. For culture, multiple tags may be present when prompts include several cues (e.g., attire + signage).
A.3.  Data Generation and Quality Control
Prompting scheme. We construct base prompts for diverse scenes (home, street, workspace, worship, nature) and attach single-factor variants that toggle one attribute at a time (gender, race, age group, religion symbol, culture cue, disability aid) while holding the rest fixed. This design supports counterfactual swaps under matched backgrounds.
Parsing and validation. We deterministically parse labels from prompt text and normalize to the target taxonomy. We block contradictory combinations (e.g., multiple single-label values), reject near-duplicates via perceptual hashing, and run image-level sanity checks for the presence/absence of specified objects or symbols.
Splits. We provide stratified {train, dev, test} splits that preserve the marginal distributions of all single-label attributes and the frequency of top culture tags. Prompt IDs and seeds allow exact regeneration.
A.4.  Recommended Uses and Tasks
DIVERSIFY enables: (1) fairness audits for attributes that are weakly encoded in pixels (religion, culture, disability); (2) evaluation of abstention and evidence use by MLLM judges vs. contrastive scorers; (3) counterfactual probes by swapping a single attribute with the scene held constant; and (4) robustness studies under prompt-template perturbations.
A.5.  Ethical Considerations and Risks
Synthetic, not personal data. Labels refer to intended depictions in generated images and must not be conflated with real identities. Stereotype risk. Some combinations may mirror societal stereotypes despite careful prompt design; we encourage per-attribute reporting and qualitative review of failures. Cultural sensitivity. Culture and religion are multi-faceted; our tags reflect visible cues and are necessarily reductive. Results should not be interpreted as normative claims about any group.
A.6.  Limitations
First, labels are prompt-derived; generators may fail to render requested cues, leading to label–image mismatch. Second, taxonomies are coarse by design; many identities and intersections are absent or underrepresented. Third, culture is multi-label and cue-driven, so counts reflect prompt content rather than self-identified culture. Finally, generator style and priors introduce artifacts that may confound downstream evaluation.
A.7.  Licensing and Availability
We release prompts, canonicalized labels, and split files under a permissive license, with scripts to regenerate images subject to the base model’s terms. [^2] License text and regeneration instructions are provided in the repository. Metadata include prompt IDs, normalized labels, and generation seeds to facilitate exact reproduction and re-scoring under alternative taxonomies.
Appendix B Dataset Contribution and Datasheet: DIVERSIFY-Professions
B.1.  Overview
We contribute a synthetic image dataset for profession-focused fairness analysis. The corpus comprises 1,200 images generated from text prompts that explicitly vary profession, gender, race, age group, religion, and culture cues. Prompts are organized as a small set of base prompts with systematic variants that control a single factor at a time, enabling controlled evaluation of representation biases in multimodal models.
Intended use.
The dataset is designed for: (i) measuring demographic skews in model outputs conditioned on profession; (ii) stress-testing captioning, retrieval, and classification models for spurious correlations; and (iii) evaluating intervention methods (e.g., debiasing prompts or reweighting) under matched visual conditions.
B.2.  Composition and Statistics
Each image is associated with structured metadata parsed from the prompt text. Table 12 shows the profession coverage; Tables 1314 summarize demographic attributes; Table 15 lists the top culture tags (multi-label).
Table 12: Profession coverage (single label; N=1,200N{=}1{,}200).
Profession Count %
Doctor 200 16.7%
Engineer 200 16.7%
Janitor 200 16.7%
Lawyer 200 16.7%
Nurse 200 16.7%
Teacher 200 16.7%
Table 13: Gender and race distributions (single label; N=1,200N{=}1{,}200).
Gender
Male 498 41.5%
Female 487 40.6%
Unspecified 215 17.9%
Race
Black 421 35.1%
White 403 33.6%
South Asian 140 11.7%
Middle Eastern 115 9.6%
East Asian 86 7.2%
Hispanic/Latino 19 1.6%
Native American 7 0.6%
Unspecified 9 0.8%
Table 14: Age groups and religion (single label; N=1,200N{=}1{,}200).
Age group
Young adult 343 28.6%
Middle-aged 374 31.2%
Senior 357 29.8%
Unspecified 126 10.5%
Religion
Christian 226 18.8%
Muslim 230 19.2%
Hindu 264 22.0%
Sikh 89 7.4%
Buddhist 55 4.6%
Jewish 10 0.8%
Taoist 6 0.5%
Unspecified 320 26.7%
Table 15: Top culture tags (multi-label; percentages do not sum to 100).
Culture tag Prompts % of NN
Unspecified 551 45.9%
South Asian 216 18.0%
Middle Eastern 203 16.9%
East Asian 137 11.4%
West African 58 4.8%
Hispanic/Latino 25 2.1%
Turkish 22 1.8%
Chinese 12 1.0%
Native American 10 0.8%
Arab/Middle Eastern 3 0.2%
Label provenance.
All labels are text-derived from prompts; no manual annotation of the rendered images was performed. For culture, multiple cues may be present; we therefore treat culture as multi-label and report unique-prompt counts per tag.
B.3.  Data Generation and Quality Control
Prompting scheme. We start from a profession-specific base prompt (e.g., “Base prompt (doctor)”) and create controlled variants (e.g., V1p1(doctor), V1p2(doctor)) to toggle a single attribute (gender, race, age group, religion, or culture cue). Parsing and validation. We deterministically extract labels from prompt text using pattern matching over explicit descriptors (male/female, Black/White/East Asian/…, young adult/middle-aged/senior, Christian/Muslim/Hindu/…) and canonicalize them to a closed taxonomy. We perform consistency checks to ensure exactly one profession per prompt and at most one label per single-label attribute.
B.4.  Recommended Uses and Tasks
This dataset supports: (1) profession-conditioned fairness probes (e.g., are nurses depicted differently by gender or race?); (2) auditing retrieval/captioning models for spurious attribute leakage; and (3) counterfactual testing by swapping a single attribute while holding the rest constant.
B.5.  Ethical Considerations and Risks
Synthetic images are not people. Demographic labels describe intended depictions in generated images; they must not be conflated with real identities. Stereotype risk. Some attribute–profession combinations may mirror societal stereotypes. We encourage reporting performance disaggregated by both profession and demographic attribute, and avoiding aggregate-only metrics. Religious and cultural sensitivity. Labels rely on visible symbols or textual cues, which can be reductionist. Downstream use should avoid normative claims about any group.
B.6.  Limitations
(i) Labels are derived from text prompts rather than human verification of image content; the generator may fail to render the intended attribute. (ii) Taxonomies are coarse (e.g., “Black”, “White”) and not exhaustive; some identities are absent or underrepresented (e.g., Hispanic/Latino 1.6%, Native American 0.6%). (iii) Culture is multi-label and cue-based, so counts reflect presence of textual/visual cues rather than self-identified culture. (iv) The visual style and biases of the underlying image generator may introduce artifacts unrelated to the prompts.
B.7.  Licensing and Availability
We release the prompts and metadata (.csv/.json) under a permissive license and provide scripts to regenerate the images where permitted by the image model’s license. [^3] Exact license text and regeneration instructions included in the project repository. To facilitate reproducibility, we include prompt IDs, attribute labels, and (where applicable) generation seeds.
B.8.  Reproducibility Checklist
  • Files: prompts.txt, metadata.csv (id, profession, gender, race, age_group, religion, culture), and a datasheet (.xlsx).
  • Counts: N=1,200N{=}1{,}200 images; 6 professions with 200 images each.
  • Determinism: Prompt $\rightarrow$ label mapping is rule-based and released; image seeds provided when licensing allows.
  • Splits: Deterministic train/dev/test split seeds provided for reproducible evaluation.
Appendix C FairJudge Prompts Templates
All prompts enforce closed label sets, require visible evidence, and allow a neutral fallback (unspecified/neutral). We parse only the structured fields during scoring; rationales are logged for audit.
Attribute prediction (single prompt shared across models).
Listing 1: Attribute prediction prompt (label-constrained with abstention).
You are given an image of a person. Classify the following using ONLY visible evidence.
If evidence is insufficient for any field, output "unspecified".
Labels:
- Gender: {male, female, unspecified}
- Race: {Black, White, East Asian, South Asian, Middle Eastern,
Latino_Hispanic, Indigenous, unspecified}
- Age: {child, young adult, middle-aged, elderly, unspecified}
- Religion: {Christian, Muslim, Hindu, Buddhist, Jewish, Sikh, Shinto,
Neutral, unspecified}
- Culture: free-text tag OR "unspecified" (e.g., British, African, East Asian,
Latin American, ...), use "unspecified" if no visible cue
- Disability: {mobility impairment, blind/low vision, deaf/hard of hearing,
dwarfism, vitiligo, unspecified}
Output STRICT JSON (no extra text), with these keys exactly:
{"Gender": "...", "Race": "...", "Age": "...",
"Religion": "...", "Culture": "...", "Disability": "..."}
Alignment judging (rubric $\rightarrow$ scalar [1,1][-1,1]).
Listing 2: Prompt–image alignment rubric (later rescaled to [-1,1]).
Task: Rate how well the image matches the description.
Description:
"<CAPTION>"
Use this 1--5 rubric:
1 = Not a match at all
2 = Significant discrepancies
3 = Several minor discrepancies
4 = A few minor discrepancies
5 = Matches exactly
Answer in STRICT JSON (no extra text):
{"Alignment analysis": "<<=100 words, evidence-based>>",
"Alignment score": "<integer 1-5>"}
Profession accuracy + alignment (optional combined prompt).
Listing 3: Combined profession + alignment judging.
You are given an image and a candidate profession. First, decide if the image shows that profession.
Then rate alignment using the 1--5 rubric above. Base both decisions ONLY on visible evidence.
Input:
Profession = "<PROFESSION>"
Caption = "a photo of a <PROFESSION>"
Output STRICT JSON:
{"Profession prediction": "<yes|no|unspecified>",
"Prediction evidence": "<<=50 words>>",
"Alignment analysis": "<<=100 words>>",
"Alignment score": "<integer 1-5>"}
Implementation notes. We keep wording identical across models (GPT-4.1, Gemini 1.5 pro 1.5 pro, LLaMA-4). At evaluation, we parse the JSON fields only and linearly rescale the 1–5 alignment rating to $\overline{s}\in[-1,1]$ for comparison with CLIP cosine; minimal template perturbations are reported in Sec. 4.4.

### v2

@@
 As a result, fairness assessments may either overstate model capability or mask failure modes that would be obvious to a careful human reviewer.
-Large-scale studies have documented representation disparities in T2I and vision–language models and have introduced diagnostic tests and scorecards for auditing them (Luccioni et al., 2023; Wan et al., 2024).
+Large-scale studies have documented representation disparities in T2I and vision–language models and have introduced diagnostic tests and scorecards for auditing them (luccioni2023stablebias; wan2024survey).
 Remedial approaches range from lightweight sampling control to small learned adaptors.
-Fair Diffusion steers the sampler toward user-specified demographic proportions without retraining the backbone (Friedrich et al., 2023); Fair Mapping inserts a linear adapter to systematically reduce demographic skew with minimal compute (Li et al., 2023).
-Generation-time prompting with chain-of-thought (CoT), as in FairCoT, leverages reasoning to mitigate bias during image synthesis (Al Sahili et al., 2024).
+Fair Diffusion steers the sampler toward user-specified demographic proportions without retraining the backbone (friedrich2023fairdiffusion); Fair Mapping inserts a linear adapter to systematically reduce demographic skew with minimal compute (li2023fairmapping).
+Generation-time prompting with chain-of-thought (CoT), as in FairCoT, leverages reasoning to mitigate bias during image synthesis (sahili2025faircot).
 While effective for mitigation, these approaches operate on the generator; they do not provide an evaluation protocol that is calibrated, evidence-grounded, and appropriately abstaining.
-On the evaluation side, contrastive similarity (e.g., CLIP) is a dominant proxy for text–image faithfulness (Radford et al., 2021; Hessel et al., 2021), and VQA-style checklists probe finer-grained grounding (Hu et al., 2023), but both can reward surface correlation and penalize uncertainty.
-Preference-trained rankers (ImageReward, PickScore) improve agreement with human choices (Xu et al., 2023; Kirstain et al., 2023) yet remain scalar scorers with limited transparency.
-Recent “LLM-as-a-judge” work reports high agreement with human raters but also reveals position and verbosity biases, and thus calls for careful protocol design and meta-evaluation (Zheng et al., 2023; Chen et al., 2024; Wang et al., 2025).
+On the evaluation side, contrastive similarity (e.g., CLIP) is a dominant proxy for text–image faithfulness (radford2021clip; hessel2021clipscore), and VQA-style checklists probe finer-grained grounding (hu2023tifa), but both can reward surface correlation and penalize uncertainty.
+Preference-trained rankers (ImageReward, PickScore) improve agreement with human choices (xu2023imagereward; kirstain2023pickapic) yet remain scalar scorers with limited transparency.
+Recent “LLM-as-a-judge” work reports high agreement with human raters but also reveals position and verbosity biases, and thus calls for careful protocol design and meta-evaluation (zheng2023llmasjudge; chen2024humansllms; wang2025mllmjudge).
 In short: we lack a lightweight, evaluation-oriented procedure that (i) reasons explicitly about visual evidence, (ii) adheres to label constraints, (iii) abstains when cues are insufficient, and (iv) applies beyond face-visible demographics to attributes like religion, culture, disability, and profession.
@@
 Text-to-image (T2I) systems inherit and sometimes amplify social biases present in their training data.
-Early large-scale analyses show systematic disparities across gender, skin tone, geo-culture, and professions, and propose diagnostic protocols and scorecards for auditing representation (Luccioni et al., 2023; Wan et al., 2024).
-Several debiasing methods have been explored post-hoc or with light adaptation: Fair Diffusion steers sampling toward user-specified demographic proportions (Friedrich et al., 2023), while Fair Mapping learns a small linear adaptor to systematically reduce demographic skew with minimal compute (Li et al., 2023).
-Concurrently, FairCoT leverages chain-of-thought (CoT) reasoning in multimodal LLMs to mitigate bias during generation-time prompting (Al Sahili et al., 2024).
+Early large-scale analyses show systematic disparities across gender, skin tone, geo-culture, and professions, and propose diagnostic protocols and scorecards for auditing representation (luccioni2023stablebias; wan2024survey).
+Several debiasing methods have been explored post-hoc or with light adaptation: Fair Diffusion steers sampling toward user-specified demographic proportions (friedrich2023fairdiffusion), while Fair Mapping learns a small linear adaptor to systematically reduce demographic skew with minimal compute (li2023fairmapping).
+Concurrently, FairCoT leverages chain-of-thought (CoT) reasoning in multimodal LLMs to mitigate bias during generation-time prompting (sahili2025faircot).
 Our work differs in purpose: rather than intervening on the generator, we study evaluation fairness by treating multimodal LLMs (MLLMs) as judges over social attributes and alignment outcomes.
@@
 Face-centric attribute benchmarks remain widely used to quantify gender, race, and age performance.
-FairFace provides a balanced dataset across seven race groups and age/gender labels (Kärkkäinen and Joo, 2021).
-For context-embedded protected-attribute cues beyond faces, the PaTA benchmark was curated alongside the DeAR debiasing method for VLMs (Seth et al., 2023).
+FairFace provides a balanced dataset across seven race groups and age/gender labels (karkkainen2021fairface).
+For context-embedded protected-attribute cues beyond faces, the PaTA benchmark was curated alongside the DeAR debiasing method for VLMs (seth2023dear).
 Tooling spans classic face-attribute models and modern vision-language recognition.
-In practice, many works rely on the open-source DeepFace toolkit to obtain age, gender, and race predictions (Serengil, 2024), and on CLIP (Radford et al., 2021) for zero-shot category attribution via text prompts.
-However, audits of CLIP-family encoders find systematic gender and racial skew whose magnitude depends on model size and, crucially, the pre-training data source/scale (Al Sahili et al., 2025).
+In practice, many works rely on the open-source DeepFace toolkit to obtain age, gender, and race predictions (serengilDeepFace), and on CLIP (radford2021clip) for zero-shot category attribution via text prompts.
+However, audits of CLIP-family encoders find systematic gender and racial skew whose magnitude depends on model size and, crucially, the pre-training data source/scale (sahili2025data).
 We extend this line by prompting MLLMs (GPT-4.1, Gemini, LLaMA-4) to directly infer social attributes, enabling coverage of non-facial categories (e.g., religion, culture, disability) with explanations.
 Text–image alignment.
-CLIP-based similarity remains a dominant automatic proxy for prompt faithfulness (Hessel et al., 2021).
-Beyond cosine-based scores, VQA-driven faithfulness frameworks such as TIFA generate question–answer checklists to probe fine-grained concept grounding (Hu et al., 2023).
-Preference-trained reward/score models (e.g., ImageReward and PickScore) leverage large-scale human comparisons to better match human judgments for T2I ranking and evaluation (Xu et al., 2023; Kirstain et al., 2023).
-Recently, X-IQE proposes a Chain-of-Thought, explanation-oriented evaluator using visual LLMs, reporting improved agreement with humans while outputting interpretable rationales (Chen et al., 2023).
+CLIP-based similarity remains a dominant automatic proxy for prompt faithfulness (hessel2021clipscore).
+Beyond cosine-based scores, VQA-driven faithfulness frameworks such as TIFA generate question–answer checklists to probe fine-grained concept grounding (hu2023tifa).
+Preference-trained reward/score models (e.g., ImageReward and PickScore) leverage large-scale human comparisons to better match human judgments for T2I ranking and evaluation (xu2023imagereward; kirstain2023pickapic).
+Recently, X-IQE proposes a Chain-of-Thought, explanation-oriented evaluator using visual LLMs, reporting improved agreement with humans while outputting interpretable rationales (chen2023xiqe).
 In our alignment study, we compare CLIP similarity against MLLM-based judging calibrated with an X-IQE–style scoring protocol to a standardized [1,1][-1,1] range.
@@
 Using LLMs as evaluators (“LLM-as-a-judge”) has become a practical alternative to costly human evaluation.
-Foundational work reports high agreement with human preferences but also identifies position, verbosity, and self-enhancement biases, motivating careful protocol design and meta-evaluation (Zheng et al., 2023).
-Subsequent studies investigate reliability, consistency, and bias in LLM judging, including on open-ended tasks (Chen et al., 2024).
-For images, recent work shows that MLLMs can serve as safety and policy judges with structured reasoning pipelines, revealing both promise and pitfalls for multimodal evaluation (Wang et al., 2025).
+Foundational work reports high agreement with human preferences but also identifies position, verbosity, and self-enhancement biases, motivating careful protocol design and meta-evaluation (zheng2023llmasjudge).
+Subsequent studies investigate reliability, consistency, and bias in LLM judging, including on open-ended tasks (chen2024humansllms).
+For images, recent work shows that MLLMs can serve as safety and policy judges with structured reasoning pipelines, revealing both promise and pitfalls for multimodal evaluation (wang2025mllmjudge).
 Our work complements this literature by: (i) framing social-attribute prediction as a judging task on face and non-face domains (e.g., culture, religion, disability), (ii) quantifying per-attribute accuracy and generalization across FairFace, PaTA, and new human-centric scenarios, and (iii) benchmarking alignment judging on IdenProf professions and FairCoT—contrasting CLIP-based scores with MLLM judges that also supply free-form rationales.
@@
 8.  Bibliographical References
-[…]
 Appendix A Dataset Contribution and Datasheet: DIVERSIFY